# Lattice-Renormalized Tunneling Models for Superconducting Qubit Materials

### v1

@@ Introduction.
In the ground state, strain coupling is strongly suppressed: tunneling defects with no intrinsic well-to-well asymmetry exhibit only second-order strain interactions. Consequently, strain is required to the break degeneracy in the stress-free state. In practice, equilibrium configurations are rarely achieved in thin films due to finite-size effects, making it essential to evaluate strain effects across a range of hydrostatic and deviatoric strains relative to the infinite-lattice geometry. In the harmonic limit, the nuclear Hamiltonian near a minimum may be expanded as $E^{(i)}/V_{0}=C_{jklm}^{(i)}\varepsilon_{jk}\varepsilon_{lm}/2$, where $C_{jklm}^{(i)}$ is the elastic stiffness tensor and $\varepsilon_{jk}$ the strain components. Under a non-zero macroscopic strain field, $\alpha^{(i)}_{jk}$, the energy of site $i$ becomes $E^{(i)}/V_{0}=\frac{1}{2}C_{jklm}^{(i)}(\varepsilon_{jk}-\alpha^{(i)}_{jk})(\varepsilon_{lm}-\alpha^{(i)}_{lm})$, which contains the first-order term:
@@
This framework demonstrates that local strain fields can strongly modulate TLS energetics, making strain engineering and control over homogeneity critical for mitigating decoherence [33]; highly pure superconducting films with low internal strain fields, will be the most prone to resonant TLS and multilevel systems (MLS). In crystalline systems, only TLS subject to nearly isotropic stress/strain fields will exhibit tunneling behavior, as tunneling is effectively quenched above a critical well-asymmetry.
Four-Level and Multi-Level Systems.
Substitutional defects in bcc Nb preserve the cubic point group symmetry. Adjacent to any substitutional impurity there are 24 degenerate tetrahedral sites. This high-site degeneracy leads to complex energy-level structures for isolated H defects which occupy those sites. The H site degeneracy, however, is affected by internal stress fields that reduce the number of degenerate or nearly-degenerate H sites [27, 25]. Indeed, anelastic-relaxation experiments on NbZr0.0045Hx only supports the existence of a population of H four-level systems (FLS) [27, 25].
[…]
@@
In reality, substitutional defects in Nb act generate MLS described by an effective 24-dimensional Hamiltonian. The exact level structure of will depend on the relative well energies which in turn depends on the local-stress state (as shown above). Like FLS, these MLS will possess a level structure essentially bounded by the hyperbolic transition-energy curves of a given TLS subspace, leaving intermediate transitions within a fixed upper limit; the lower limit is bound by zero. Consequently, even substantial lattice strains cannot fully eliminate such MLS from contributing as resonant or subresonant states under stress fields that would typically quench two-level behavior. These MLS maintain transition energies within the qubit frequency range over a broader span of internal strain fields; however, due to the increased site degeneracy, this behavior persists only when the local strain tensor is nearly isotropic. For example, Zr-H defects in the dilute limit have been experimentally shown to support at most a four-level structure, far below the theoretical upper bound of 24 [27, 25]. While heavier lattice masses and strain fields suppress TLS activity, MLS remain robust over a wider range of conditions, making them a critical consideration for understanding decoherence in superconducting qubits.
Conclusions.
Standard approaches in the superconducting qubit community fail to capture essential features of two-level and multilevel defects arising from degenerate minima in the nuclear Hamiltonian. Symmetrized structures significantly overestimate tunnel splittings and are not thermodynamically stable, particularly for configurations with large coincidence energies. We address this limitation by introducing composite phonon coordinates that account for lattice displacements between degenerate wells, enabling a lattice-renormalized formalism for configurational tunneling systems.
@@ Conclusions.
Our calculations for O-H and O-D tunneling defects in bcc Nb demonstrate that a 4D Hamiltonian provides a lower bound on experimental tunnel splittings 0.064 meV (0.0048 meV) compared to 0.19 meV (0.021 meV) for oxygen-trapped H (D) TLS in bcc Nb, while a 3D model without $Q$ bounds them from above 0.57 meV (0.040 meV). Additionally, $J_{\mathrm{H}}$ computed for self-trapped hydrogen is comparable to $J_{\mathrm{D}}$ computed for oxygen-trapped hydrogen. This finding supports the conclusion that almost all H is either trapped by defect sites or precipitates to the $\varepsilon$-phase at mK temperature; as otherwise, the measured specific heat anomaly should extend below that observed in [23, 29]. Extending to a 5D Hamiltonian for a FLS enhances the splitting by 27%, primarily due to bending of the tunneling path enabled by phonon coupling. These results highlight the critical role of anharmonic coupling to lattice phonons near the barrier maximum, suggesting that further refinement of the model will improve agreement with experiment.
@@
Our formalism extends naturally to multilevel systems, which remain active under strain conditions that suppress TLS-qubit interactions and thus represent an important source of decoherence in superconducting qubits. By linking tunneling dynamics to phonon-mediated strain interactions, we show that local stress fields, via elastic dipole coupling, strongly modulate defect energetics while the composite coordinate $Q$ introduces additional channels for energy exchange beyond conventional strain coupling. We further note that the TLS density for amorphous materials is on the order of $0.1$ eV-1 m-3 [34]. The H TLS density may be estimated as $\rho_{\mathrm{H}}Z(E)\approx 2\rho_{H}/\pi\varepsilon_{0}$ [23]. Using $\varepsilon_{0}=4.5$ meV and $\rho_{\mathrm{H}}=0.43$ nm-3 ($c=1/128$), we find the TLS density to be approximately 60 eV-1 nm-3. Correspondingly, even a dilute density of interstitial H introduces several orders of magnitude more TLS than the same volume of a glassy material. These insights underscore the need for strain engineering and homogeneity control in superconducting films and provide a predictive framework for designing materials that minimize TLS-related loss in superconducting quantum circuits.
@@ Appendix A Derivation of the reduced Hamiltonian
$V(\{\vec{r}_{i}\})\approx\sum_{k}H_{k}(\vec{r}_{H})\left[V_{k}^{(0)}+\frac{1}{2}\sum_{ij}\sum_{\mu\nu}\left(r_{i,\mu}-r_{i,\mu}^{\,eq,k}\right)\Phi^{(k)}_{i\mu,j\nu}\left(r_{j,\nu}-r_{j,\nu}^{\,eq,k}\right)\right]+V_{\text{anharm}}(\{\vec{r}_{i}\}),$ (3)
where $\Phi^{(k)}$ is the Hessian in basin $k$, and $V_{\text{anharm}}$ collects higher-order and inter-basin couplings. (Here $i$ and $j$ are atom indices, $k$ indexes each degenerate H site, and $\mu$ and $\nu$ are the cartesian indices for the atomic coordinates $\vec{r}_{i}$ and $\vec{r}_{j}$.) Inserting Eq. (3) into Eq. (2) gives the harmonic Hamiltonian used in the main text.
The mass dependence may be eliminated by using the standard transformation to phonon coordinates:
@@
$\hat{H}_{n}=\sum_{i}\frac{-\hbar^{2}}{2}{\nabla^{\prime}}_{i}^{2}+\sum_{ik}\frac{1}{2}\omega_{ik}^{2}q_{ik}^{2}H_{k}(\vec{q}_{H})+V_{anharm}(\{q_{i}\})$ (5)
Where $\nabla^{\prime}$ indicates the mass-scaled $\nabla$ operator, $w_{ij}$ and $q_{ij}$ represent the frequency and mode amplitude of phonon mode $i$ in the harmonic well $k$. For clarity, we do no explicitly distinguish the longitudinal and transverse modes. We also make the approximation that $\mathbf{q}=\vec{q}_{H}=m_{H}^{1/2}\vec{r}_{H}$ which is justified as the phonon eigenvector of the H modes in bcc Nb posses essentially no contribution from Nb atoms. Because the kinetic energy is invariant under orthonormal changes of basis in mass-weighted coordinates, the Laplacian transforms as $\sum_{i}(-\hbar^{2}/2m_{i})\nabla_{i}^{2}\to\sum_{\alpha}(-\hbar^{2}/2)\,\partial^{2}/\partial Q_{\alpha}^{2}$. This justifies the use of the mass-scaled operator $\nabla^{\prime}$.

### v2

@@ IV.5 Acoustic Phonons and Strain Interactions
In the ground state, strain coupling is strongly suppressed: tunneling defects with no intrinsic well-to-well asymmetry exhibit only second-order strain interactions. Consequently, strain is required to break the degeneracy in the stress-free state. In practice, equilibrium configurations are rarely achieved in thin films due to finite-size effects, making it essential to evaluate strain effects across a range of hydrostatic and deviatoric strains relative to the infinite-lattice geometry. In the harmonic limit, the nuclear Hamiltonian near a minimum may be expanded as $E^{(i)}/V_{0}=C_{jklm}^{(i)}\varepsilon_{jk}\varepsilon_{lm}/2$, where $C_{jklm}^{(i)}$ is the elastic stiffness tensor and $\varepsilon_{jk}$ the strain components. Under a non-zero macroscopic strain field, $\alpha^{(i)}_{jk}$, the energy of site $i$ becomes $E^{(i)}/V_{0}=\frac{1}{2}C_{jklm}^{(i)}(\varepsilon_{jk}-\alpha^{(i)}_{jk})(\varepsilon_{lm}-\alpha^{(i)}_{lm})$, which contains the first-order term:
@@
This framework demonstrates that local strain fields can strongly modulate TLS energetics, making strain engineering and control over homogeneity critical for mitigating decoherence [33]; highly pure superconducting films with low internal strain fields, will be the most prone to resonant TLS and multilevel systems (MLS). In crystalline systems, only TLS subject to nearly isotropic or likewise symmetry-preserving stress/strain fields will exhibit tunneling behavior, as tunneling is effectively quenched above a critical well-to-well asymmetry.
@@
IV.6 Four-Level and Multi-Level Systems
Substitutional defects in bcc Nb preserve its cubic point group symmetry. Adjacent to any substitutional impurity there are 24 degenerate tetrahedral sites. This high-site degeneracy leads to complex energy-level structures for isolated H defects which occupy those sites. The H site degeneracy, however, is affected by internal stress fields that reduce the number of degenerate or nearly-degenerate H sites [27, 25]. Indeed, anelastic-relaxation experiments on NbZr0.0045Hx only supports the existence of a population of H four-level systems (FLS) [27, 25].
[…]
@@ IV.6 Four-Level and Multi-Level Systems
In reality, substitutional defects in Nb generate MLS described by an effective 24-dimensional Hamiltonian. The exact level structure will depend on the relative well energies which in turn depends on the local-stress state (as shown above) and short-range defect-defect interactions. Like FLS, these MLS will possess a level structure essentially bounded by the hyperbolic transition-energy curves of a given TLS subspace, leaving intermediate transitions within a fixed upper limit; the lower limit is bound by zero. Consequently, even substantial lattice strains cannot fully eliminate such MLS from contributing as resonant or subresonant states under stress fields that would typically quench two-level behavior. These MLS maintain transition energies within the qubit frequency range over a broader span of internal strain fields; however, due to the increased site degeneracy, this behavior persists only when the local strain tensor is nearly isotropic. For example, Zr-H defects in the dilute limit have been experimentally shown to support at most a four-level structure, far below the theoretical upper bound of 24 [27, 25]. While heavier lattice masses and strain fields suppress TLS activity, MLS remain robust over a wider range of conditions, making them a critical consideration for understanding decoherence in superconducting qubits.
IV.7 Tunnel splittings from machine-learned potentials
A recent work employed a machine-learned potential (MLP) to compute tunnel splittings for O-trapped H in bcc Nb, reporting $J=0.275$ meV for the low-energy trapping site corresponding to our site 7a, and $J=0.414$ meV for a higher energy site (equivalent to our site 5a). These values were obtained using a static lattice in a symmetric configuration following a procedure closely related to that described in Sec. C. We note that the identification of the most stable O-H configuration and the associated relaxation protocol match those employed, although not documented, in an earlier draft of the present work [34], which predates the results of Ref. [35].
The $J$ reported in Ref. [35] is most directly comparable to our static-lattice DFT result in Table 4, where we obtain $J_{3D}=0.57$ meV for for O-trapped H at site 7a. The MLP used in Ref. [35] is reported to achieve a mean absolute error in the computed total energy per atom of less than 1 meV/atom, which is below the typical $\sim$5 meV/atom threshold typically considered acceptable for high-quality MLPs. However, no errors in the predicted atomic positions were reported, despite the fact that even picometer deviations in H coordinates and lattice positions can strongly impact the computed tunnel splitting.
This comparison emphasizes an important limitation in interpreting tunnel splittings obtained from MLPs: Even for chemically simple systems, such as O-trapped H in bcc Nb, uncertainties intrinsic to the MLP parametrization introduce an additional source of error beyond those inherent to the underlying DFT functional. While a well-trained MLP may produce splittings in reasonable agreement with experiment, such agreement alone does not validate either the parametrization strategy or the theoretical formalism and may in fact be accidental. Apparent improvements over DFT can arise from error cancellation within the interatomic potential and should be acknowledged. Similar forms of cancellation, and hence accidental agreement, are documented and occur in high-throughput low-fidelity DFT studies [36, 37].
Our results show that a static-lattice treatment at the DFT level systematically overestimates the experimental tunnel splitting, yielding $J_{\mathrm{3D}}$ values larger than measured. In contrast, inclusion of coupling to the composite mode $Q$ suppresses the splitting substantially, producing $J_{\mathrm{4D}}$ values which bound the experimental $J$ from below by an equal margin. We show in a forthcoming work that the inclusion of the lattice mode which passes through the transition state alleviates most of the remaining error in the computation of $J$. We would expect the same qualitative behavior, i.e., an over-estimation of $J$ in the static‑lattice approximation and a large reduction when coupling to the $Q$ mode is included, using the MLP of Ref. [35].
V Conclusions
@@
Standard approaches in the superconducting qubit community fail to capture essential features of two-level and multilevel defects arising from degenerate minima in the nuclear Hamiltonian. Symmetrized structures significantly overestimate tunnel splittings and are not thermodynamically stable, particularly for configurations with large coincidence energies. We address this limitation by introducing composite phonon coordinates that account for lattice displacements between degenerate wells, enabling a lattice-renormalized formalism for configurational tunneling systems. Our calculations for O-H and O-D tunneling defects in bcc Nb demonstrate that a 4D Hamiltonian provides a lower bound on experimental tunnel splittings 0.064 meV (0.0048 meV) compared to 0.19 meV (0.021 meV) for oxygen-trapped H (D) TLS in bcc Nb, while a 3D model without $Q$ bounds them from above 0.57 meV (0.040 meV). Additionally, $J_{\mathrm{H}}$ computed for self-trapped hydrogen is comparable to $J_{\mathrm{D}}$ computed for oxygen-trapped hydrogen. This finding supports the conclusion that almost all H is either trapped by defect sites or precipitates to the $\varepsilon$-phase at mK temperature; as otherwise, the measured specific heat anomaly should extend below that observed in [23, 30]. Extending to a 5D Hamiltonian for a FLS enhances the splitting by 27%, primarily due to bending of the tunneling path enabled by additional phonon coupling. These results highlight the critical role of anharmonic coupling to lattice phonons near the barrier maximum, suggesting that further refinement of the model will improve agreement with experiment.
By explicitly computing the formation energy of H tetrahedral sites adjacent to each trapping defect, we show that O, Ti, and Zr-trapped H exhibit unambiguously favored sites. Each of the thermodynamically favored sites possess at least one adjacent, degenerate site which validates the assumption of active tunneling systems in H-doped Nb with secondary trapping defects [22, 23, 27]. In particular, O-trapped H hosts only 2 adjacent, degenerate sites supporting the two-level model proposed in [23]. In contrast, Ti- and Zr-trapped H possess 24 adjacent, degenerate sites that support possible multi-level system behavior of these defects [27, 25]. The preference for these sites are well-explained by simple elastic and quantum mechanical bonding arguments.
@@
Our formalism extends naturally to multilevel systems, which remain active under strain conditions that suppress TLS-qubit interactions and thus represent an important source of decoherence in superconducting qubits. By linking tunneling dynamics to phonon-mediated strain interactions, we show that local stress fields, via elastic dipole coupling, strongly modulate defect energetics while the composite coordinate $Q$ introduces additional channels for energy exchange beyond conventional strain coupling. We further note that the TLS density for amorphous materials is on the order of $0.1$ eV-1 nm-3 [38]. The H TLS density may be estimated as $\rho_{\mathrm{H}}Z(E)\approx 2\rho_{H}/\pi\varepsilon_{0}$ [23]. Using $\varepsilon_{0}=4.5$ meV and $\rho_{\mathrm{H}}=0.43$ atoms nm-3 ($c=1/128$), we find the TLS density to be approximately 60 eV-1 nm-3. Correspondingly, even a dilute density of interstitial H introduces several orders of magnitude more TLS than the same volume of a glassy material. These insights underscore the need for strain engineering and homogeneity control in superconducting films and provide a predictive framework for designing materials that minimize TLS-related loss in superconducting quantum circuits.
@@ Appendix A Derivation of the reduced Hamiltonian
$V(\{\vec{r}_{i}\})=\sum_{k}H_{k}(\vec{r}_{H})\left[V_{k}^{(0)}+\frac{1}{2}\sum_{ij}\sum_{\mu\nu}\left(r_{i,\mu}-r_{i,\mu}^{\,eq,k}\right)\Phi^{(k)}_{i\mu,j\nu}\left(r_{j,\nu}-r_{j,\nu}^{\,eq,k}\right)\right]+V_{\text{anharm}}(\{\vec{r}_{i}\}),$ (4)
where $\Phi^{(k)}$ is the Hessian in basin $k$, and $V_{\text{anharm}}$ collects higher-order and inter-basin couplings. (Here $i$ and $j$ are atom indices, $k$ indexes each degenerate H site, and $\mu$ and $\nu$ are the cartesian indices for the atomic coordinates $\vec{r}_{i}$ and $\vec{r}_{j}$.)
The mass dependence of Eq. 3 may be eliminated by using the standard transformation to phonon coordinates:
@@
$\hat{H}_{n}=\sum_{i}\frac{-\hbar^{2}}{2}{\nabla^{\prime}}_{i}^{2}+\sum_{ik}\frac{1}{2}\omega_{ik}^{2}q_{ik}^{2}H_{k}(\vec{q}_{\mathrm{H}})+V_{anharm}(\{q_{i}\})$ (6)
Where $\nabla^{\prime}$ indicates the mass-scaled $\nabla$ operator, $w_{ij}$ and $q_{ij}$ represent the frequency and mode amplitude of phonon mode $i$ in the harmonic well $k$. For clarity, we do no explicitly distinguish the longitudinal and transverse modes. We also make the approximation that $\mathbf{q}=\vec{q}_{\mathrm{H}}=m_{\mathrm{H}}^{1/2}\vec{r}_{\mathrm{H}}$ which is justified as the phonon eigenvector of the H modes in bcc Nb posses essentially no contribution from Nb atoms. Because the kinetic energy is invariant under orthonormal changes of basis in mass-weighted coordinates, the Laplacian transforms as $\sum_{i}(-\hbar^{2}/2m_{i})\nabla_{i}^{2}\to\sum_{\alpha}(-\hbar^{2}/2)\,\partial^{2}/\partial Q_{\alpha}^{2}$. This justifies the use of the mass-scaled operator $\nabla^{\prime}$.
@@
Appendix B Nb Phonons
To evaluate the vibrational coupling between Nb, O, and H ions in bcc Nb with an O-H defect, we computed the $\Gamma$-point phonon frequencies and eigenvectors for the $c=1/128$ structure with an O-H defect using the Phonopy code [41]. The computed H mode frequencies are 29.5, 41.4, and 41.5 THz, whereas the O mode frequencies are 9.97, 11.4, and 19.0 THz. The phonon eigenvectors were converted to relative displacements by scaling them by the mass matrix $\mathbf{M}^{-1/2}=m_{i}^{-1/2}\delta_{ij}$, where $m_{i}$ is the mass of ion $i$. The computed H displacement component of each H mode exceeded 0.9999, while for each O mode the component of the displacement exceeded 0.99. These results strongly justify our treatment of the H modes as fully decoupled from all other atoms.
Fig. A1 shows the distribution of mode frequencies and the projection of the composite mode $\mathbf{\hat{Q}}$ onto the phonon eigenvectors $\mathbf{\hat{e}}_{i}$. A single dominant, low-frequency mode is present with a projected value of 0.50. The remaining modes with non-zero projections are broadly distributed across the phonon frequency spectrum—consistent with the localized displacements which transform the lattice between degenerate configurations.
Figure A1: (top) $\Gamma$-point phonon frequencies for the OH defect structure ($c=1/128$) in bcc Nb. (bottom) Projection of the composite coordinate unit vector $\mathbf{\hat{Q}}$ onto each phonon eigenvector $\mathbf{\hat{e}}_{i}$. Oxygen and hydrogen modes are excluded in both panels.
Appendix C Coincidence Structures
In Ref. [28], a theory is introduced in which light tunneling particles are treated adiabatically with respect to lattice. The lattice relaxes under an effective force that can be derived from the Hellman-Feynman theorem:
$\mathbf{F_{n}}=\int d\mathbf{r}\left|\psi(\mathbf{r})\right|^{2}\mathbf{f}_{n}(\mathbf{r})\,$, (10)
where $\psi(\mathbf{r})$ is the wavefunction of the light particle, $\mathbf{f}_{n}(\mathbf{r})$ is the force on lattice ion $n$ due to the light particle located at position $\mathbf{r}$, and $\mathbf{F}_{n}$ is the resulting effective force on ion $n$. This formulation is conveniently adapted for computing coincidence structures.
When the probability density is approximated as a superposition of two delta functions centered at the most probable positions of the light particle, the effective force on the lattice reduces to:
$\mathbf{F}_{n}=\left|\psi(\mathbf{r}_{1})\right|^{2}\mathbf{f}_{n}(\mathbf{r}_{1})+\left|\psi(\mathbf{r}_{2})\right|^{2}\mathbf{f}_{n}(\mathbf{r}_{2})\,.$ (11)
An equivalent equation can be derived for the stress tensor $\sigma_{ij}$. For H or D interstitials in bcc Nb, the coincidence configuration corresponds to equal wavefunction population in each potential well. We emphasize that relaxing the lattice using Eq. 10 is equivalent to a constrained relaxation of the lattice subject to an equal population constraint. Eq. 10 was implemented in combination with the BFGS optimizer in the Atomic Simulation Environment [42] (ASE) to obtain fully relaxed coincidence structures. The coincidence energy, $E_{c}$, is defined as the relative energy difference between the coincidence structure and the self-trapped configuration. As shown in Table 3, the coincidence energy determined obtained using Eq. 10 is approximately 10% lower than the effective coincidence energy $E_{c}^{\prime}$.